OpenStudy (anonymous):

find the general solution for sinx=1-2cos^2x. I have already found that x=7pi/6, 11pi/6 and pi/2; I just need help finding the general solution. I can never understand how to find the general solution ;_; thankss!! (last question i promise lol)

OpenStudy (anonymous):

@ganeshie8 wanna help me again :x or anyone really

OpenStudy (kc_kennylau):

\[\Large\begin{array}{rcl} \sin x&=&1-2\cos^2x\\ \sin x&=&1-2(1-\sin^2x)\\ \sin x&=&1-2+2\sin^2x\\ 2\sin^2x-\sin x-1&=&0\\ (2\sin x+1)(\sin x-1)&=&0\\ 2\sin x+1=0&\text{or}&\sin x-1=0\\ \sin x=-0.5&\text{or}&\sin x=1\\ x=\frac{11\pi+2m\pi}6\text{ or }x=\frac{7\pi+2n\pi}6&\text{or}&x=\frac{\pi+2p\pi}2 \end{array}\]

OpenStudy (kc_kennylau):

@ganeshie8 lolz u faster :P

OpenStudy (anonymous):

thanks @kc_kennylau, but can you explain the method in which you arrived at the general solution? and what is the difference between the m and n?

OpenStudy (kc_kennylau):

just giving 'em different names :)

OpenStudy (kc_kennylau):

wait sorry it should be: \[\Large x=\frac{7\pi}6+2m\pi\text{ or }x=\frac{11\pi}6+2n\pi\text{ or }x=\frac\pi2+2p\pi\]

OpenStudy (kc_kennylau):

because \(\sin\)'s period is \(2\pi\) :)

OpenStudy (anonymous):

ohh so its as easy as adding 2pi? because I knew that rule but I didn't know if I could apply it here. btw how do you type it in the cool equation form thing in different lines? I try doing that but it only allows me to enter one continuous line?

OpenStudy (kc_kennylau):

do you know how to type latex instead of just clicking the equation button? :)

OpenStudy (kc_kennylau):

like really typing out the code

OpenStudy (anonymous):

nope. oh well. thanks for the help anyways!